# Perturbative and Non-linear Analyses of Gravitational Turbulence in Spacetimes with Stable Light Rings

### v1

@@ IV.2 Four-dimensional Model
Without loss of generality, we will now set $b=1$. The challenging part about extracting the behavior of these norms from numerical simulations is that, due to numerical errors, we only have access to the lower order multipoles, roughly, $\ell\lesssim 16$. In particular, for the values of $\alpha,R$ that we choose for numerical purposes, $\alpha=0.166$ and $R=0.229$ (chosen to agree with [67], which make it simpler for the potential well to be numerically resolvable after compactifying the domain), the damping time of the $\ell=2$ mode is only $\tau_{20}/R\sim 10^{3}$ (see Fig. 2). If we excite initially the $\ell=2$ mode, we will only see the norm growth on timescales shorter than this decay timescale, $t\ll\tau_{20}$. From the $2$-dimensional model, we may expect that the norm grows as $\epsilon^{4}t^{2}$, as shown in Eq. 28. The high power of the amplitude $\epsilon$ in this scaling makes it challenging to observe the norm growth for small initial data, even for the higher order norms.
@@ V.1 Two-Dimensional Model
As done in the previous Section, let us start with the lower-dimensional model, i.e., the nonlinear cubic wave equation in the circle. We evolve numerically the system of Eq. (19) with a pseudospectral strategy. We use the ApproxFun.jl package [75] to compute the nonlinear term in real space, and then evolve the Fourier representation of the equation in time using an explicit, fourth order Runge–Kutta method. After each step, we dealias the latter half of the Fourier modes to filter out contributions that can numerically contaminate the lower-frequency modes [76]. A convergence test for our numerical solutions is shown in App. D.
We prepare the evolution by exciting the two modes $n=10,11$, both with amplitude $\epsilon=20$ (i.e., far away from the expected regime of validity of perturbation theory), and evolve the system until (long times) $T\gtrsim\epsilon^{-2}$. In Fig. 4 we present the evolution of the resulting spectrum, where we can see the development of an inertial range, characterized by a polynomial law, which can be interpreted as a direct energy cascade towards higher frequencies.
Figure 4: Field spectrum $|\phi_{n}|$ at different wavenumber $n$ for large initial data with amplitude $\epsilon=20$, initially excited in the modes $n=10,11$. The different colors indicate the field spectrum at different times, as indicated by the colorbar. The spectrum at the final time of the simulation, $t=5000\approx\epsilon^{2}$ is shown as a wider, red line. It agrees with the predicted Kolmogorov–Zakharov scaling $n^{-3/2}$ (dashed black line). The resolution of this simulation is $N=1024$ modes.
@@
There are special solutions to the kinetic wave equation which are stationary, i.e., $\dot{n}_{k}=0$, under the assumption that there is dissipation acting at low and high frequencies, and the system is being injected energy at some scale [77, 78] (which is not the case for the model we are interested). In this scenario, there exist solutions of the form $n_{k}\sim k^{-\nu}$, in the regime bounded by the dissipation scales, which is known as the inertial regime. These spectra are known as Kolmogorov–Zakharov (KZ) spectra, and there are many techniques devoted to compute the power of these spectra $\nu$ precisely. For four wave equations, the direct energy cascade is associated to a power $\nu=3/2$ [85]. In Fig. 4 we are comparing the late-time solution with this prediction and found good agreement.
Equipped with these expectations, we can now address the question about the end state of the nonlinear wave equation in this two-dimensional model, for small initial data. If there is no dissipation, a direct energy cascade towards higher energies ensues, ultimately spreading energy across all possible modes. In the asymptotic state, $t\to\infty$ all modes are excited with negligible amplitude (since the initial amount of energy in the system was finite), and the solution loses regularity in the asymptotic time, as evidenced by the fact that the higher order norms diverge as $t\to\infty$. However, in most physical situations we expect the system to dissipate energy beyond some scale $k\geq k_{\rm Diss}$. The development of a KZ-like inertial range requires of continuous injection of energy, precisely to compensate this dissipation. In the absence of energy injection, the end state would just be $\Phi=0$, i.e., the trivial solution. We emphasize that other nonlinear equations (for example, those for which the global existence of solutions is not guaranteed) or initial data very far from the perturbative regime, where this analysis based on weak wave turbulence does not hold, might exhibit different phenomenology at late times, such as, e.g., time-periodic solutions.
@@ V.2 Nonlinear Waves on the Sphere
where we have re-scaled time as $T=\sqrt{f(R)}t$. We examine the nonlinear wave equation in this geometry, i.e., $\Box_{g_{R}}\Phi=\kappa\Phi^{3}$, where now $\Phi:\mathcal{S}^{2}\times\mathbf{R}\to\mathbf{R}$ is a real field. Expanding in spherical harmonics, and assuming axisymmetry, we can write $\Phi=\sum_{\ell}\phi_{\ell}(T)\mathcal{Y}_{\ell,0}(\theta)$, so that the nonlinear wave equation reads
@@
is a purely oscillatory term, which combines the homogeneous modes and the non-linearly excited (cubic) modes. Note that we write $\Omega^{\sigma_{1}\sigma_{2}}_{\ell_{0}}=\Omega^{\sigma_{1}\sigma_{2}}_{(\ell_{0},\ell_{0},\ell_{0})}$. Therefore, for $\ell\neq\ell_{0}$ (e.g., for $\ell=3\ell_{0}$) the solution is purely oscillatory, and it excites both homogeneous and cubic modes with an amplitude which is cubic in the linear amplitude $\mathcal{A}_{\ell_{0}}$. The $\ell=\ell_{0}$ mode, however, features a secular linear growth in time, as well as sub-leading, oscillatory pieces. The results are the same for a general set of initially excited modes, $\tilde{\Lambda}$. Generically we have that $\phi_{\ell}^{(3)}\sim t\sin(\omega_{\ell}T)$ if $\ell\in\tilde{\Lambda}$, and otherwise, $\phi_{\ell}^{(3)}\sim\cos(\omega_{\ell}T)$ whenever $\ell\notin\tilde{\Lambda}$. This implies that all energy norms grow at the same rate, independently of the order of the norm.
@@
and similarly for the other two possible combinations. Therefore, as $\ell$ grows more modes become almost exactly resonant. This dispersion relation is also the dispersion relation of linear waves on the Möbius open strip. Intuitively, this can be seen as arising from the fact that axisymmetric waves on the sphere are waves in the closed interval $[0,\pi]$, subject to the periodicity relation $\Phi(\theta)=\Phi(\theta+2\pi)$. A simple way to achieve this is by requiring the anti-periodicity condition $\Phi(\theta)=-\Phi(\theta+\pi)$, which under an appropriate rescaling is the condition naturally satisfied by sections of the fiber bundle associated to the Möbius strip double cover of the circle.
@@
Figure 9: The power spectrum $|\phi_{\ell}|$ at different times (as indicated by the colorbar), for the same configuration presented in Fig. 8. The latest time available is represented with a thick-dark red line. At high frequencies, the spectrum is consistent with a Kolmogorov–Zakharov-type scaling $\phi_{\ell}\sim\ell^{-3/2}$ (black dashed line), as in the two dimensional model.
@@ VI Discussion and Future Outlook
Several horizonless black hole mimickers and higher-dimensional spacetimes are known to admit stable light rings, which can efficiently trap high-frequency radiation, including gravitational waves. Consequently, small perturbations in such spacetimes decay slowly [44, 51]. While it has been conjectured that this slow decay could trigger a nonlinear instability, the existing numerical evidence remains inconclusive. In particular, no direct numerical confirmation of such an instability has been observed in specific models without additional factors, such as a breakdown of axisymmetry [48], Recent numerical work [67] has provided evidence that a stable LR can induce a turbulent cascade capable of transferring energy to higher frequencies. In this work, we extend this analysis by further investigating the dynamics of this cascade in the perturbative regime. Our results indicate that the sole presence of a stable LR does not necessarily lead to instability. In particular, for the nonlinear cubic wave equation considered here, we argue that the trivial solution $\Phi=0$ remains stable under small perturbations.
@@
A key question is whether this turbulent cascade could still lead to a physical instability over sufficiently long timescales. To address this, we have studied the long-term dynamics of a dimensionally reduced system, which describes waves propagating exactly at the LR. This model effectively captures the dynamics of $-\ell$ modes, and our numerical simulations reveal hints of a dual cascade–where energy flows both towards higher frequencies (direct cascade) and towards lower frequencies (inverse cascade). When artificial dissipation is introduced at high frequencies, an inertial range emerges, characterized by a Kolmogorov–Zakharov spectrum. However, it is important to recognize that this model may not accurately describe the high-frequency behavior. In particular, a family of long-lived radial overtones exists for large angular numbers, supported within a finite-thickness spherical shell centered around the LR. This suggests that conclusions from the reduced model should be treated cautiously, and further research is needed.
@@
If the considered system with a stable light ring were to be unstable, two possible outcomes could be expected: either the formation of a black hole at the LR or the emergence of time-periodic solutions. However, since radiation is only partially trapped within a finite-width shell, the formation of a black hole starting from arbitrarily small initial data appears inconceivable. This expectation aligns with the arguments presented in [67]. Unlike the case of pure AdS, where nonlinear wave interactions lead to the focusing of energy at a single point [52], in this type of spacetimes, the energy remains distributed across an extended region. Additionally, we find no evidence for the formation of time-periodic solutions, aside from a turbulent Kolmogorov–Zakharov spectra. However, for these characteristic spectra to be maintained, both (i) high-frequency dissipation and (ii) a continuous energy injection mechanism are required. In the absence of sustained energy input, they must eventually decay. Our results, therefore, suggest that, after a sufficiently long time, the radiation either escapes the light ring or disperses into high-frequency modes with negligible amplitude. Based on this analysis, nonlinear waves propagating in spacetimes with stable light rings appear dynamically stable within the framework considered here.
@@
Several aspects remain unexplored in this work. If a black hole mimicker also possesses a horizon, in addition to a stable light ring, dissipation is enhanced, reinforcing the arguments against an instability. However, as noted in Ref. [94], topological considerations imply that in four-dimensional asymptotically flat spacetimes, the presence of a horizon and a stable light ring necessitates a second stable light ring, introducing additional complexity. The interplay between multiple stable light rings warrants further investigation. Additionally, our analysis is based on a scalar field model, which, while capturing some essential nonlinear dynamics, does not account for full gravitational interactions. Since trapped radiation at the light ring can, in principle, modify the light ring’s properties, the backreaction effects in a fully relativistic setting could introduce additional dynamical features. Lastly, rotation can amplify radiation, either through partial absorption at the surface of the black hole mimicker or via the presence of ergoregions [95]. Studying these effects in a well-posed matter model remains an important next step in confirming the conclusions drawn here.
@@ Acknowledgements.
We are very grateful to Frans Pretorius for inspiring the two-dimensional model we studied and for his feedback on all the stages of this work. We also thank Gabriele Benomio, Vitor Cardoso, and Nils Siemonsen for helpful discussions and comments. J.R-Y. is grateful to the Princeton Gravity Initiative for hospitality during the early stages of development of this work. J.R-Y. acknowledges support by VILLUM Foundation (grant no. VIL37766) and the DNRF Chair program (grant no. DNRF162) by the Danish National Research Foundation. The Center of Gravity is a Center of Excellence funded by the Danish National Research Foundation under grant No. 184. A.C.-A. acknowledges support from the DOE through Los Alamos National Laboratory (LANL) Directed Research and Development, grant 20240748PRD1, as well as by the Center for Nonlinear Studies. This work is authorized for unlimited release under LA-UR-25-21674. The Tycho supercomputer hosted at the SCIENCE HPC center at the University of Copenhagen was used for supporting this work.
@@ Appendix A The WKB Approximation
In this section we will briefly review the application of the Werners–Kramers–Brillouin (WKB) approximation to find the long-lived modes of a system. The approximation in this case corresponds to a perturbative expansion in $1/\ell$, and thus is accurate for high wavenumbers. However, it has also been shown to be very accurate [73, 45] in the presence of long-lived modes, i.e., in order to capture modes of the system with very large quality factors.
@@
This condition is just the Bohr–Sommerfeld quantization rule
@@
We can now compute the leading order scaling of the frequencies with the angular number $\ell$. Indeed, by expanding around the stable light ring $r\sim R$, and keeping only the leading order term, $V\sim\ell^{2}f(R)/R^{2}$, the Bohr–Sommerfeld condition yields
@@
Appendix B Breit–Wigner Resonances
Chandrasekhar and Ferrari [96, 97] noticed that compact stars close to the Buchdahl limit have long-lived modes. Usually we think of QNMs as poles of the scattering matrix, but an alternative viewpoint is to think them as Breit–Wigner resonances in the scattering amplitude. By using this idea, they designed an algorithm which is very efficient at finding the frequencies of these longest lived modes.
We start by integrate outwards, ensuring that the solution is regular at the origin. At large distances, the outgoing solution takes the form $\phi_{\ell}\sim A\cos(\omega r)+B\sin(\omega r)$. Near the frequency of a long-lived mode, the scattering amplitude has the following behaviour:

### v2

@@ IV.2 Four-dimensional model
Without loss of generality, we will now set $b=1$. The challenging part about extracting the behavior of these norms from numerical simulations is that, due to numerical errors, we only have access to the lower order multipoles, roughly, $\ell\lesssim 16$. In particular, for the values of $\alpha,R$ that we choose for numerical purposes, $\alpha=0.166$ and $R=0.229$ (chosen to agree with [68], which make it simpler for the potential well to be numerically resolvable after compactifying the domain), the damping time of the $\ell=2$ mode is only $\tau_{20}/R\sim 10^{3}$ (see Fig. 2). If we excite initially the $\ell=2$ mode, we will only see the norm growth on timescales shorter than this decay timescale, $t\ll\tau_{20}$. From the $2$-dimensional model, we may expect that the norm grows as $\epsilon^{4}t^{2}$, as shown in Eq. (28). The high power of the amplitude $\epsilon$ in this scaling makes it challenging to observe the norm growth for small initial data, even for the higher order norms.
@@ V.1 Two-dimensional Model
As done in the previous section, let us start with the lower-dimensional model, i.e., the nonlinear cubic wave equation in the circle. We evolve numerically the system of Eq. (19) with a pseudospectral strategy. We use the ApproxFun.jl package [76] to compute the nonlinear term in real space, and then evolve the Fourier representation of the equation in time using an explicit, fourth order Runge-Kutta method. After each step, we dealias the latter half of the Fourier modes to filter out contributions that can numerically contaminate the lower-frequency modes [77]. A convergence test for our numerical solutions is shown in Appendix D.
We prepare the evolution by exciting the two modes $n=10,11$, both with amplitude $\epsilon=20$ (i.e., far away from the expected regime of validity of perturbation theory), and evolve the system until (long times) $T\gtrsim\epsilon^{-2}$. In Fig. 4 we present the evolution of the resulting spectrum, where we can see the development of an inertial range, characterized by a polynomial law, which can be interpreted as a direct energy cascade toward higher frequencies.
Figure 4: Field spectrum $|\phi_{n}|$ at different wave number $n$ for large initial data with amplitude $\epsilon=20$, initially excited in the modes $n=10,11$. The different colors indicate the field spectrum at different times, as indicated by the color bar. The spectrum at the final time of the simulation, $t=5000\approx\epsilon^{2}$ is shown as a wider, red line. It agrees with the predicted Kolmogorov-Zakharov scaling $n^{-3/2}$ (dashed black line). The resolution of this simulation is $N=1024$ modes.
@@
There are special solutions to the kinetic wave equation which are stationary, i.e., $\dot{n}_{k}=0$, under the assumption that there is dissipation acting at low and high frequencies, and the system is being injected energy at some scale [78, 79] (which is not the case for the model we are interested). In this scenario, there exist solutions of the form $n_{k}\sim k^{-\nu}$, in the regime bounded by the dissipation scales, which is known as the inertial regime. These spectra are known as Kolmogorov-Zakharov (KZ) spectra, and there are many techniques devoted to compute the power of these spectra $\nu$ precisely. For four wave equations, the direct energy cascade is associated to a power $\nu=3/2$ [86]. In Fig. 4 we are comparing the late-time solution with this prediction and found good agreement.
Equipped with these expectations, we can now address the question about the end state of the nonlinear wave equation in this two-dimensional model, for small initial data. If there is no dissipation, a direct energy cascade toward higher energies ensues, ultimately spreading energy across all possible modes. In the asymptotic state, $t\to\infty$ all modes are excited with negligible amplitude (since the initial amount of energy in the system was finite), and the solution loses regularity in the asymptotic time, as evidenced by the fact that the higher order norms diverge as $t\to\infty$. However, in most physical situations we expect the system to dissipate energy beyond some scale $k\geq k_{\rm Diss}$. The development of a KZ-like inertial range requires of continuous injection of energy, precisely to compensate this dissipation. In the absence of energy injection, the end state would just be $\Phi=0$, i.e., the trivial solution. We emphasize that other nonlinear equations (for example, those for which the global existence of solutions is not guaranteed) or initial data very far from the perturbative regime, where this analysis based on weak wave turbulence does not hold, might exhibit different phenomenology at late times, such as, e.g., time-periodic solutions.
@@ V.2 Nonlinear waves on the sphere
where we have rescaled time as $T=\sqrt{f(R)}t$. We examine the nonlinear wave equation in this geometry, i.e., $\Box_{g_{R}}\Phi=\kappa\Phi^{3}$, where now $\Phi:\mathcal{S}^{2}\times\mathbf{R}\to\mathbf{R}$ is a real field. Expanding in spherical harmonics, and assuming axisymmetry, we can write $\Phi=\sum_{\ell}\phi_{\ell}(T)\mathcal{Y}_{\ell,0}(\theta)$, so that the nonlinear wave equation reads
@@
is a purely oscillatory term, which combines the homogeneous modes and the nonlinearly excited (cubic) modes. Note that we write $\Omega^{\sigma_{1}\sigma_{2}}_{\ell_{0}}=\Omega^{\sigma_{1}\sigma_{2}}_{(\ell_{0},\ell_{0},\ell_{0})}$. Therefore, for $\ell\neq\ell_{0}$ (e.g., for $\ell=3\ell_{0}$) the solution is purely oscillatory, and it excites both homogeneous and cubic modes with an amplitude which is cubic in the linear amplitude $\mathcal{A}_{\ell_{0}}$. The $\ell=\ell_{0}$ mode, however, features a secular linear growth in time, as well as subleading, oscillatory pieces. The results are the same for a general set of initially excited modes, $\tilde{\Lambda}$. Generically we have that $\phi_{\ell}^{(3)}\sim t\sin(\omega_{\ell}T)$ if $\ell\in\tilde{\Lambda}$, and otherwise, $\phi_{\ell}^{(3)}\sim\cos(\omega_{\ell}T)$ whenever $\ell\notin\tilde{\Lambda}$. This implies that all energy norms grow at the same rate, independently of the order of the norm.
@@
and similarly for the other two possible combinations. Therefore, as $\ell$ grows more modes become almost exactly resonant. This dispersion relation is also the dispersion relation of linear waves on the Möbius open strip. Intuitively, this can be seen as arising from the fact that axisymmetric waves on the sphere are waves in the closed interval $[0,\pi]$, subject to the periodicity relation $\Phi(\theta)=\Phi(\theta+2\pi)$. A simple way to achieve this is by requiring the antiperiodicity condition $\Phi(\theta)=-\Phi(\theta+\pi)$, which under an appropriate rescaling is the condition naturally satisfied by sections of the fiber bundle associated to the Möbius strip double cover of the circle.
@@
Figure 9: The power spectrum $|\phi_{\ell}|$ at different times (as indicated by the color bar), for the same configuration presented in Fig. 8. The latest time available is represented with a thick-dark red line. At high frequencies, the spectrum is consistent with a Kolmogorov-Zakharov-type scaling $\phi_{\ell}\sim\ell^{-3/2}$ (black dashed line), as in the two dimensional model.
@@ VI Discussion and Future Outlook
Several horizonless black hole mimickers and higher-dimensional spacetimes are known to admit stable light rings, which can efficiently trap high-frequency radiation, including gravitational waves. Consequently, small perturbations in such spacetimes decay slowly [44, 51]. While it has been conjectured that this slow decay could trigger a nonlinear instability, the existing numerical and analytical evidence remains inconclusive.
Recent numerical work [68] has provided evidence that a stable LR can induce a turbulent cascade capable of transferring energy to higher frequencies. In this work, we extend this analysis by further investigating the dynamics of this cascade in the perturbative regime. Our results indicate that the sole presence of a stable LR does not necessarily lead to instability. In particular, for the nonlinear cubic wave equation considered here, we argue that the trivial solution $\Phi=0$ remains stable under small perturbations.
@@
A key question is whether this turbulent cascade could still lead to a physical instability over sufficiently long timescales. To address this, we have studied the long-term dynamics of a dimensionally reduced system, which describes waves propagating exactly at the LR. This model effectively captures the dynamics of $-\ell$ modes, and our numerical simulations reveal hints of a dual cascade—where energy flows both toward higher frequencies (direct cascade) and toward lower frequencies (inverse cascade). When artificial dissipation is introduced at high frequencies, an inertial range emerges, characterized by a Kolmogorov-Zakharov spectrum. However, it is important to recognize that this model may not accurately describe the high-frequency behavior. In particular, a family of long-lived radial overtones exists for large angular numbers, supported within a finite-thickness spherical shell centered around the light rings. This suggests that conclusions from the reduced model should be treated cautiously, and further research is needed.
@@
If the considered system with a stable light ring, were to be unstable, and the nonlinear backreaction of the scalar field on the geometry were included, at least the following possible outcomes could be expected: the formation of a black hole at the LR; the emergence of time-periodic solutions (e.g., quasimode solutions exhibiting an uniform decay rate slower than inverse logarithmically in time); or the evolution toward a state without a stable LR altogether (as, e.g., conjectured in Ref. [48] for spinning boson stars but recently reconsidered in Ref. [95]). Regarding the first scenario, since radiation is only partially trapped within a finite-width shell, the formation of a black hole starting from arbitrarily small initial data appears inconceivable. This expectation aligns with the arguments presented in [68]. Unlike the case of pure AdS, where nonlinear wave interactions lead to the focusing of energy at a single point [52], in this type of spacetimes, the energy remains distributed across an extended region. Additionally, we find no evidence for the formation of time-periodic solutions, aside from a turbulent Kolmogorov–Zakharov spectra. However, for these characteristic spectra to be maintained, both (i) high-frequency dissipation and (ii) a continuous energy injection mechanism are required. In the absence of sustained energy input, they must eventually decay. Our results, therefore, suggest that, after a sufficiently long time, the radiation either escapes the light ring or disperses into high-frequency modes with negligible amplitude. Based on this analysis, nonlinear waves propagating in spacetimes with stable light rings appear dynamically stable within the framework considered here.
@@
Several aspects remain unexplored in this work. If a black hole mimicker also possesses a horizon, in addition to a stable light ring, dissipation is enhanced, reinforcing the arguments against an instability. However, as noted in Ref. [96], topological considerations imply that in four-dimensional asymptotically flat spacetimes, the presence of a horizon and a stable light ring necessitates a second unstable light ring, introducing additional complexity. The interplay between multiple light rings warrants further investigation. Additionally, our analysis is based on a scalar field model, which, while capturing some essential nonlinear dynamics, does not account for full gravitational interactions. Since trapped radiation at the light ring can, in principle, modify the light ring’s properties, the backreaction effects in a fully relativistic setting could introduce additional dynamical features that the nonlinear behavior arising solely from the self-interaction of the scalar field does not capture. Lastly, rotation can amplify radiation, either through partial absorption at the surface of the black hole mimicker or via the presence of ergoregions [97]. Studying these effects in a well-posed matter model remains an important next step in confirming the conclusions drawn here.
@@ Acknowledgements.
We are very grateful to Frans Pretorius for inspiring the two-dimensional model we studied and for his feedback on all the stages of this work. We also thank Gabriele Benomio, Vitor Cardoso, and Nils Siemonsen for helpful discussions and comments. J.R-Y. is grateful to the Princeton Gravity Initiative for hospitality during the early stages of development of this work. J.R-Y. acknowledges support by VILLUM Foundation (grant no. VIL37766) and the DNRF Chair program (grant no. DNRF162) by the Danish National Research Foundation. The Center of Gravity is a Center of Excellence funded by the Danish National Research Foundation under grant No. 184. A.C.-A. acknowledges support from the DOE through Los Alamos National Laboratory (LANL) Directed Research and Development, Grant No. 20240748PRD1, as well as by the Center for Nonlinear Studies. This work is authorized for unlimited release under LA-UR-25-21674. The Tycho supercomputer hosted at the SCIENCE HPC center at the University of Copenhagen was used for supporting this work.
@@ Appendix A The WKB Approximation
In this section we will briefly review the application of the Wentzel-Kramers-Brillouin (WKB) approximation to find the long-lived modes of a system. The approximation in this case corresponds to a perturbative expansion in $1/\ell$, and thus is accurate for high wave numbers. However, it has also been shown to be very accurate [74, 45] in the presence of long-lived modes, i.e., in order to capture modes of the system with very large quality factors.
@@
This condition is just the Bohr-Sommerfeld quantization rule
@@
We can now compute the leading order scaling of the frequencies with the angular number $\ell$. Indeed, by expanding around the stable light ring $r\sim R$, and keeping only the leading order term, $V\sim\ell^{2}f(R)/R^{2}$, the Bohr-Sommerfeld condition yields
@@
Appendix B Breit-Wigner Resonances
Chandrasekhar and Ferrari [98, 99] noticed that compact stars close to the Buchdahl limit have long-lived modes. Usually we think of QNMs as poles of the scattering matrix, but an alternative viewpoint is to think them as Breit-Wigner resonances in the scattering amplitude. By using this idea, they designed an algorithm which is very efficient at finding the frequencies of these longest lived modes.
We start by integrate outward, ensuring that the solution is regular at the origin. At large distances, the outgoing solution takes the form $\phi_{\ell}\sim A\cos(\omega r)+B\sin(\omega r)$. Near the frequency of a long-lived mode, the scattering amplitude has the following behavior: